OpenStudy (anonymous):

the linearization of the function f= sqrt(x) at x =9 is?

OpenStudy (anonymous):

L(x)= f(a)+f'(a)(x-a) just plug 9 in for a.

OpenStudy (anonymous):

L(9)=3-1/6(x-9) is the answer i believe

OpenStudy (experimentx):

Oo ... i think @chemENGINEER is right ... i have been doing lot of above stuffs ...

OpenStudy (anonymous):

the correct ans according to the practice test i am looking at is. y= x/6 + 3/2, and the approximation of sqrt(8.5) by the linearization is 2.9167

OpenStudy (shayaan_mustafa):

now put x=8.5 is that you are asking for?

OpenStudy (experimentx):

looks like i really screwed up above, i took the point as 0 myself and took the function as (1+x)^(1/2) (x)^(1/2) = 3 + (x-9)/6 same as above = 3/2 + x/6

OpenStudy (experimentx):

you must have misread chemengineer .. (x-9) was at numerator